OpenStudy (anonymous):

An aquarium sells cubic fish tanks of different sizes. The length of the small size fish tank is 4 feet. The dimensions of the jumbo size fish tank are double the dimensions of the small size fish tank. Which expression can be used to find the ratio of the volume of the small size fish tank to the volume of the jumbo size fish tank? (8/4)^2 (4/8)^3 (8/4)^3 (4/8)^2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

for similar figures, if dimensions are in ratio \(\large \frac{m}{n}\) , then, areas will be in ratio \(\large (\frac{m}{n})^2\) Volumes will be in ratio \(\large (\frac{m}{n})^3\)

OpenStudy (anonymous):

B?

ganeshie8 (ganeshie8):

Yup ! always look for ^3 for volumes, ^2 for area